1.use Cramer's rule to solve the equation below (10 marks)


y-z =2

3x+2y+z=4

5x+4y=1


1
Expert's answer
2022-04-05T17:17:11-0400

"\Delta=\begin{vmatrix}
 0& 1&-1 \\
 3& 2&1\\
5&4&0
\end{vmatrix}=0+(-1)^3(-5)-(-1)^4(12-10)=3"

"\Delta_{x}=\begin{vmatrix}
 2& 1&-1 \\
 4& 2&1\\
1&4&0
\end{vmatrix}=(-1)^22(-4)+(-1)^3(-1)+(-1)^4(-1)(16-2)=-8+1-14=-21"

"\Delta_{y}=\begin{vmatrix}
 0& 2&-1 \\
 3& 4&1\\
5&1&0
\end{vmatrix}=(-1)^32(-5)+(-1)^4(-1)(3-20)=27"

"\Delta_{z}=\begin{vmatrix}
 0& 1&2 \\
 3& 2&4\\
5&4&1
\end{vmatrix}=(-1)^3(3-20)+(-1)^42(12-10)=21"

x=-21/3=-7

y=27/3=9

z=21/3=7